# MistralBSM: Leveraging Mistral-7B for Vehicular Networks Misbehavior Detection

### v1

@@ III-B Transformers and LLMs
Figure 2: Transformer architecture
III-C Low rank adaptation
In traditional fine-tuning, the model’s weights are directly modified to adapt the model to the new considered task. While fine-tuning is a crucial approach ensuring transfer learning, the process become computationally daunting as the number of parameter increases. Notably, fine-tuning large language models poses significant challenges in terms of memory requirements and computational resources.
Low Rank Adaptation (LoRA), is a technique that was proposed in [34] to alleviate the memory footprint and speed up the process of large models’ fine-tuning. This is achieved through reducing the number of trainable parameters by decomposing the changes on the pre-trained weight matrices into low rank matrices. The process suppose that after fine-tuning the new weights matrices will have the form $W=W+\Delta W$ where W is frozen during training and $\Delta W$ is decomposed into two lower rank matrices A and B. Being of lower rank, the number of trainable parameters in these matrices is smaller than the original $\Delta W$.
@@ IV System model
We present an approach centered around the utilization of Mistral-7B for detecting vehicle misbehavior. Our method involves reframing the misbehavior detection problem as a sequence classification task. In this framework, we deploy a model based on Mistral-7B, which is quantized and implemented on a roadside unit (RSU). Through quantization, the memory requirements for running Mistral-7B are significantly reduced, enabling its deployment on edge equipment at the RSUs. These RSUs possess the necessary capabilities to collect BSMs exchanged within their sensing ranges and execute the model to identify potential misbehavior.
The collected messages undergo the prepossessing steps presented in Figure 3-a, where the irrelevant features are deleted, the messages are grouped by sender pseudonym and then transformed into BSM-Windows of size $n$ for each user. Following the prepossessing steps, the formed time windows are transformed into textual prompts necessary for the language model inference. The prompt design steps transform the multivariate time window to a text, as described in Figure 3-b. The obtained prompt is then fed to a Tokenizer and then Mistral-7B-based classifier for the final labeling. The last step following the behavior assessment is the reporting phase, where the RSU communicates the information’s on the misbehaving vehicle to a trust authority (TA) located in the cloud. The TA is then responsible for verifying the report, and feeding it, if necessary, to the cloud LLM, which can complement the report by conducting deeper analysis and validation, and incorporating additional information from a broader dataset stored in the cloud. Subsequently, the TA takes the appropriate response that could include revoking the vehicle’s access to the network.
It is important to highlight that our methodology ensures the examination of both contextual and content aspects within the messages at the edge LLM. Contextual dimensions are considered during the grouping of messages into time windows, facilitating the analysis of temporal patterns and relationships. Simultaneously, the content aspect is considered by a detailed inspection of the message contents.
@@ V Experiments and Results
In this section, we delve into the details of the conducted experiments. We start by presenting the preprocessing steps we applied on the database used, namely the VeReMi dataset. Thereafter, we define the performances metrics used to evaluate the performance of our model. After this, we present the two major sets of experiments, binary and multiclass classification. For each set, we investigate two possible methods of leveraging Mistral-7B: one involving fine-tuning and the other without. Our aim is to assess whether we can leverage Mistral-7B with minimal modifications, maximizing its utility without extensive adjustments. From there, we compare its performance against two other LLMs, LLAMA2-7B and roBERTa. Lastly, we analyze the effect of varying window sizes on the inference time to identify the necessary requirements and potential solutions for deploying the models.
@@ V-A Extended VeReMi dataset
VeReMi [36] is a Vehicular Reference of Misbehavior dataset that was created to evaluate the performance of misbehavior detection systems in the VANETs. It was generated using the Framework For Misbehavior Detection (F2MD), which is an extension of the vehicular network simulator VEINS, and a subsection of the Luxembourg SUMO Traffic (LuST) scenario covering 1.61 km² with a peak density of 67.4 vehicles per km². To experiment and assess our approach, we used the extended version of VeReMi[37] that offers more diverse attack types including DoS, Random Positioning Attacks, Constant Positioning Attacks, Disruptive Attacks, Constant Speed Attacks, Delayed Messages Attack, Data Replay Attacks, Traffic Sybil Attacks, and Eventual Stop Attacks that were simulated under varying traffic conditions specifically from 07 to 09 am and from 2 to 4 pm. The dataset is structured into trace files containing the exchange of BSMs by individual vehicles. During each attack simulation, vehicle behaviors are stored in separate files encompassing various fields including: type, receive time, sending time, sender ID, sender Pseudo, message ID, position, position noise, speed, speed noise, acceleration, acceleration noise, heading, and heading noise.
@@
Preprocessing of the raw files involved extracting all messages, concatenating them, and selecting only type-3 messages, which describe the received BSMs. Duplicate entries resulting from the interception of the same message by multiple vehicles were eliminated. Then fields describing noise attributes such as position noise, heading noise, acceleration noise and speed noise were dropped. In order to minimize memory usage during fine-tuning and inference, adjustments were made to the structure of certain fields. The heading, position, acceleration, and speed fields, originally consisting of lists of three values each representing the x, y, and z coordinates, underwent transformation. For speed and acceleration, the norm was calculated, condensing the data into single values. Additionally, the position and heading fields were split into two distinct fields: position-x and position-y, and heading-x and heading-y, respectively. Given that the z coordinate remained consistently null across all entries, it was removed. Lastly, to further reduce memory requirements, all numerical values were rounded to three decimal points. Following this, each message was labeled based on its sender ID, therefore, every message issued by a misbehaving car was labeled as attack. This classification is derived from the title of the corresponding trace file, referencing the attack-code correspondence outlined in Table I. Once the dataset was labeled, messages for each sender ID were grouped, ordered by the receive time and segmented into time windows.
@@
The final preprocessing step, is the textualization of the time windows, which could be achieved either by considering each row as a separate sentence having the feature values preceded by the corresponding field name, or column wise like described in Fig. 3-b. As the former results in large prompt size due to repeated field names, we resorted to the column-wise approach.
For our experiments, We selected a subset of attacks from the VeReMi dataset, and we processed them to extract time windows. In order to be able to select what is the best number of packets inside a time window, we created different datasets with 4 sizes, 10, 20, 50 and 100. Table I and Figure 4 give an overview of the selected attacks and class distribution after processing.
@@
TABLE I: Mapping attack type to codes[37]
@@ V-B Performance Metrics
Recall=True PositivesTrue Positives+False $\text{Recall}=\frac{\text{True Positives}}{\text{True Positives}+\text{False Negatives}}$ (2)
@@
Precision=True PositivesTrue Positives+False $\text{Precision}=\frac{\text{True Positives}}{\text{True Positives}+\text{False Positives}}$ (3)
@@
Accuracy=True Positives+True NegativesTotal $\text{Accuracy}=\frac{\text{True Positives}+\text{True Negatives}}{\text{TotalInstances}}$ (5)
V-C Binary classification
@@ V-C1 Fine-tuned Mistral
We chose to perform the fine-tuning on the self-attention modules and the weights of the classification head to maximize performance while minimizing parameter count, as recommended in [34]. Furthermore, the $r$ and the LoRA alpha in the Table II refer to the dimension of the low-rank matrices and the scaling factor which calibrates the weights update magnitude during the training, respectively. Figure 5 presents the effect of the hyperparameter r on the accuracy when window size equals 10. The figure shows that the best accuracy is obtained for an r value of to 2, whereas it decreases significantly for higher values. This justifies the choice indicated in Table II.
@@
TABLE II: Hyperparameter settings for fine-tuning with QLoRA
@@ V-C2 Mistral-7B as feature extractor
TABLE IV: Binary classification head hyperparameters
@@
TABLE V: Binary MistralFeatureNet results
V-C3 Other LLMs for binary classification
To further validate the model choice, we fine-tuned roBERTa and LLAMA2-7B under similar conditions. We employed similar qLoRA parameter described in TableII. With LLAMA2-7B we obtained good results with only 1000 instance per class, whereas roBERTa required a minimum of 3000 per class to converge. Figure 6 summarizes the different models performance across the time windows. The immediate observation is the superiority of fine-tuned Mistral-7B over all the other models on most metrics, notably the accuracy and the F1 score. Conversely, RoBERTa, yields the worst results for all metrics. This is due to it being the only model with a predefined maximum sequence length of 512, i.e., roBERTa cannot process token sequences of size surpassing 512. The most common solution to circumvent this limitation is to truncate the sequence to fit the model’s requirements. However, most of our prompts surpass this value. The effect of the truncation is moderate for window size 10, however it becomes more prominent as the size of the prompts grows. Another contributing factor to this behavior is the textualization step described in Section V-A, in which each column is considered as a sentence, so the truncation step results in eliminating whole features from the sequence.
Finally, Figure 6 suggests that 100 is the most suitable window size for the fine-tuned Mistral-7B. In contrast, 50 is the most suitable for LLAMA2-7B and MistralFeatureNet as they exhibit a deterioration when the window size is increased further.
@@ V-C3 Other LLMs for binary classification
V-D Multiclass classification
For multi-class classification, the goal is to use Mistral-7B to detect and correctly classify attacks among the 9 classes described in Table I. We employ the identical reasoning as in the binary categorization. We fine-tune Mistral-7B, and then we compare its results against, MistralFeatureNet for the multi-class classification, LLAMA2-7B and roBERTa.
@@ V-D1 Fine-tuned Mistral-7B
Figure 8 shows the performance results for different window sizes. As with binary classification, the hypothesis that the model performs better with a larger window is validated. The refined model obtained an F1 score of 0.96 in a 9-class classification task with just 1,000 instances per class. The confusion matrix in Figure 7 also proves the potential of the model, which correctly discerns most classes.
@@ V-D2 Mistral for feature extraction
TABLE VI: Classification network hyperparameters
@@
TABLE VII: Multi-class MistralFeatureNet results
V-D3 Other LLMs for multiclass classification
The performance of LLAMA2-7B and roBERTa on the nine-class classification task is illustrated in Figure 8. During training, LLAMA2-7B was configured similarly to Mistral-7B, with each class containing 1000 instances. In contrast, roBERTa required a minimum of 4000 instances per class. The results highlight the superiority of large Transformers’ over smaller ones. It is observed that LLAMA-7B achieves a similar accuracy to Mistral-7B, both around 0.95. Conversely, roBERTa consistently exhibits the least beneficial outcomes. Finally, we find that larger models, particularly fine-tuned Mistral-7B, exhibit superior processing capabilities for complex data messages and consistently outperform smaller LLMs such as roBERTa in the multiple-class classification task. Moreover, they achieve this capacity with a reduced number of samples, a crucial advantage when fewer vehicle communication data samples are available at the edge, which can occur due to fluctuations in traffic.
@@ V-E Inference time and deployment considerations
Due to the size of the Mistral-7B model, deploying it requires a balance between multiple hyperparameters to avoid adding overhead to a process that requires speed and accuracy. One method to mitigate memory demands is through quantization. In our experiments we used 4-bit quantization, which reduces significantly the memory required to load the model to 4GB.
In addition to memory requirements, it is also very important to study the inference latency. It is evident that the size of the prompt is one of the main factors affecting inference time. While we demonstrate that the best window size for binary and multiclass classification is 100 using fine-tuned Mistral-7B, opting for these values might introduce inference latency which is limitating to the global performance of the MDS. Figure 9, illustrates the evolution of the inference time with regard to the window size for Mistral-7B and LLAMA2-7B. We observe similar trends for both models, where the inference time increases notably as the window-size increases. This patern may pose a constraint on the use of LLMs in this context. Importantly, these measurements were conducted on A100 40GB GPUs and depending on the hardware, actual values may vary from the reported figures.
However, these models are, designed to seamlessly process prompts with different sizes without having to modify the architecture or apply additional preprocessing, which guarantees flexibility in terms of the window-size. Therefore, based on the computational resources available, the size can be adapted to enhance performance without notable accuracy reduction. the figures have shown that Mistral-7B. maintains good performance even with smaller window sizes.

### v2

@@ IV System model
We present an approach centered around the utilization of Mistral-7B for detecting vehicle misbehavior. Our method involves reframing the misbehavior detection problem as a sequence classification task. In this framework, we deploy a model based on Mistral-7B, which is quantized and implemented on a RSU. Through quantization, the memory requirements for running Mistral-7B are significantly reduced, enabling its deployment on edge equipment at the RSUs. These RSUs possess the necessary capabilities to collect BSMs exchanged within their sensing ranges and execute the model to identify potential misbehavior.
The process of classifying vehicle behaviors consists of three steps, as shown in Fig.3. The first step involves data preprocessing, detailed in submodule (a) of Fig.3. In this stage, features are extracted from detected BSMs in the RSU’s local database, following the data preparation steps detailed in V-A. Thereafter, the BSM messages are grouped by sender pseudonym and transformed into BSM-Windows of size $n$ for each vehicle. The second step, represented by the prompt design module in Fig.3, involves converting multivariate time windows into a textual format. The text structure is as follows:
reported values: $Feature_{1}$:[value_1,value_2,,value_n][value\_1,value\_2,...,value\_n], $Feature2$: [value_1,value_2,,value_n][value\_1,value\_2,...,value\_n], ..., $Feature_{M}$: [value_1,value_2,,value_n][value\_1,value\_2,...,value\_n]
The third and final step is classification, handled by the MistralBSM model in submodule (c) of Fig.3. This module consists of a Tokenizer, Mistral-7B layers for prompt processing, and a classification head that assigns a behavior label. Following the classification, the RSU reports the information about misbehaving vehicles is communicated to a Trust Authority (TA) located in the cloud.
The TA is, thus, responsible for verifying the reported misbehavior. If further analysis is needed, the TA forwards the report to a cloud-based LLM, which enhances the report by conducting deeper analysis and incorporating additional data from a broader dataset stored in the cloud. For instance, the cloud LLM can access historical data on the vehicle’s behavior, as well as sensor data from roadside infrastructure such as LIDAR or radar, allowing for a more thorough assessment of whether the vehicle was transmitting accurate information. This deeper analysis is implemented at the cloud level rather than at the edge for privacy reasons, since only the TA has access to the vehicle’s real identity. By performing this task in the cloud, sensitive information about the driver or vehicle remains protected and is only accessed when necessary, such as in cases where misbehavior needs to be verified before revoking network access. Subsequently, the TA takes the appropriate response that could include revoking the vehicle’s access to the network. In this paper, we assume the existence of such a cloud-based model and focus our study on the edge-LLM operating at the RSU. In fact, MistralBSM’s methodology ensures the examination of both contextual and content aspects within the messages at the edge LLM. Contextual dimensions are considered during the grouping of messages into time windows, facilitating the analysis of temporal patterns and relationships. Simultaneously, the content aspect is considered by a detailed inspection of the message contents.
@@ V Experiments and Results
In this section, we delve into the details of the conducted experiments. We start by presenting the preprocessing steps we applied on the database used, namely the VeReMi dataset. Thereafter, we define the performance metrics used to evaluate our model. Thereafter, we present the two major sets of experiments: binary and multiclass classification. For each set, we investigate two possible methods of leveraging Mistral-7B: one involving fine-tuning and the other without. We aim to investigate the possibility of leveraging Mistral-7B with minimal modifications, maximizing its utility without extensive adjustments. From there, we compare its performance against two other LLMs, LLAMA2-7B and RoBERTa. Lastly, we analyze the effect of varying window sizes on the inference time to identify the requirements and potential solutions for deploying the models.
@@ V-A Extended VeReMi Dataset
VeReMi [43] is a Vehicular Reference of Misbehavior dataset that was created to evaluate the performance of misbehavior detection systems in the VANETs. It was generated using the Framework For Misbehavior Detection (F2MD), which is an extension of the vehicular network simulator VEINS, and a subsection of the Luxembourg SUMO Traffic (LuST) scenario covering 1.61 km² with a peak density of 67.4 vehicles per km². To experiment and assess our approach, we used the extended version of VeReMi[44] that offers more diverse attack types, including DoS, Random Positioning Attacks, Constant Positioning Attacks, Disruptive Attacks, Constant Speed Attacks, Delayed Messages Attacks, Data Replay Attacks, Traffic Sybil Attacks, and Eventual Stop Attacks that were simulated under varying traffic conditions, specifically from 07 to 09 am and from 2 to 4 pm. The dataset is structured into trace files containing the exchange of BSMs by individual vehicles. During each attack simulation, vehicle behaviors are stored in separate files encompassing various fields including: type, receive time, sending time, sender ID, sender Pseudo, message ID, position, position noise, speed, speed noise, acceleration, acceleration noise, heading, and heading noise.
Preprocessing of the raw files involved extracting all messages, concatenating them, and selecting only type-3 messages, which describe the received BSMs. Duplicate entries resulting from the interception of the same message by multiple vehicles were eliminated. Then fields describing noise attributes such as position noise, heading noise, acceleration noise, and speed noise were dropped.
@@
In order to optimize memory usage during fine-tuning and inference, we made adjustments to some fields. For instance, we calculated the Euclidean norm for speed and acceleration fields in order to compress the 3-value features into a single value. Additionally, the position and heading fields were split into two distinct fields: position-x and position-y, and heading-x and heading-y, respectively. Given that the z coordinate remained consistently null across all entries, it was removed. Lastly, to further reduce memory requirements, all numerical values were rounded to three decimal points. Thereafter, each message is labeled based on its sender ID; therefore, every message issued by a misbehaving car was labeled as attack. This classification is derived from the title of the corresponding trace file, referencing the attack-code correspondence outlined in Table I. Once the dataset was labeled, messages for each sender ID were grouped, ordered by the receive time, and segmented into time windows. Importantly, we removed the sender ID, sender Pseudo, message ID features to avoid label leaking during training.
@@
The final preprocessing step is the textualization of the time windows, which could be achieved either by considering each row as a separate sentence having the feature values preceded by the corresponding field name or column-wise, like described in the submodule (b) in Fig. 3. As the former results in large prompt size due to repeated field names, we resorted to the column-wise approach.
For our experiments, we consider the two prominent types of misbehaviors in V2X communications: faulty-hardware-like behavior and purely malicious attacks. Therefore, we select Constant Position, Constant Speed, and Eventual Stop as the former type and Data Replay, Dos, Dos Random, and Dos Random Sybil as the malicious attacks from the extended VeRiMi dataset, and we preprocess them to extract BSM sequences as time windows.
TABLE I: Mapping attack type to codes
@@
In order to be able to select what is the best number of packets inside a time window, we created different datasets with 4 sizes, 10, 20, 50 and 100. Table I and Fig.4 give an overview of the selected attacks and class distribution after processing.
@@ V-B Performance Metrics
$\text{Recall}=\frac{\text{True Positives}}{\text{True Positives}+\text{False Negatives}}$ (2)
@@
$\text{Precision}=\frac{\text{True Positives}}{\text{True Positives}+\text{False Positives}}$ (3)
@@
$\text{Accuracy}=\frac{\text{True Positives}+\text{True Negatives}}{\text{Total Instances}}$ (5)
V-C Misbehavior Detection
@@ V-C1 Fine-tuned Mistral
We chose to perform the fine-tuning on the self-attention modules and the weights of the classification head to maximize performance while minimizing parameter count, as recommended in [41]. Furthermore, the $r$ and the LoRA alpha in the Table II refer to the dimension of the low-rank matrices and the scaling factor which calibrates the weights update magnitude during the training, respectively. Fig. 5 presents the effect of the hyperparameter r on the accuracy when window size equals 10. The figure shows that the best accuracy is obtained for an r value of to 2, whereas it decreases significantly for higher values. This justifies the choice indicated in Table II.
@@
TABLE II: Hyperparameter settings for fine-tuning with QLoRA
@@
TABLE III: Binary Mistral-7B results
@@ V-C2 Mistral-7B as Feature Extractor
TABLE V: Binary MistralFeatureNet results
@@
V-C3 Other LLMs for Attack Detection
To further validate the model choice, we fine-tuned RoBERTa and LLAMA2-7B under similar conditions. We employed similar qLoRA parameter described in TableII. With LLAMA2-7B we obtained good results with only 1000 instance per class, whereas RoBERTa required a minimum of 3000 per class to converge. Fig. 6 summarizes the different models performance across the time windows. The immediate observation is the superiority of fine-tuned Mistral-7B over all the other models on most metrics, notably the accuracy and the F1 score. Conversely, RoBERTa, yields the worst results for all metrics. This is due to it being the only model with a predefined maximum sequence length of 512, i.e., RoBERTa cannot process token sequences of size surpassing 512. The most common solution to circumvent this limitation is to truncate the sequence to fit the model’s requirements. However, most of our prompts surpass this value. The effect of the truncation is moderate for window size 10, however it becomes more prominent as the size of the prompts grows. Another contributing factor to this behavior is the textualization step described in Section V-A, in which each column is considered as a sentence, so the truncation step results in eliminating whole features from the sequence.
Finally, Fig. 6 suggests that 100 is the most suitable window size for the fine-tuned Mistral-7B. In contrast, 50 is the most suitable for LLAMA2-7B and MistralFeatureNet as they exhibit a deterioration when the window size is increased further.
@@ V-C3 Other LLMs for Attack Detection
V-D Misbehavior Classification
For misbehavior classification, the goal is to use Mistral-7B to detect and correctly classify attacks among the 9 classes described in Table I. We employ the identical reasoning as in the binary categorization. We fine-tune Mistral-7B, and then we compare its results against, MistralFeatureNet for the multi-class classification, LLAMA2-7B and RoBERTa.
@@ V-D1 Fine-tuned Mistral-7B
Fig. 8 shows the performance results for different window sizes. As with binary classification, the hypothesis that the model performs better with a larger window is validated. The refined model obtained an F1 score of 0.96 in a 9-class classification task with just 1,000 instances per class. The confusion matrix in Fig. 7 also proves the potential of the model, which correctly discerns most classes.
@@ V-D2 Mistral For Feature Extraction
TABLE VII: Multi-class MistralFeatureNet results
@@
V-D3 Other LLMs for Attack Classification
The performance of LLAMA2-7B and RoBERTa on the nine-class classification task is illustrated in Fig. 8. During training, LLAMA2-7B was configured similarly to Mistral-7B, with each class containing 1000 instances. In contrast, RoBERTa required a minimum of 4000 instances per class. The results highlight the superiority of large Transformers’ over smaller ones. It is observed that LLAMA-7B achieves a similar accuracy to Mistral-7B, both around 0.95. Conversely, RoBERTa consistently exhibits the least beneficial outcomes. Finally, we find that larger models, particularly fine-tuned Mistral-7B, exhibit superior processing capabilities for complex data messages and consistently outperform smaller LLMs such as RoBERTa in the multiple-class classification task. Moreover, they achieve this capacity with a reduced number of samples, a crucial advantage when fewer vehicle communication data samples are available at the edge, which can occur due to fluctuations in traffic.
@@ V-E Inference Time and Deployment Considerations
[…]
On the other hand, inference time is influenced by many factors, including the type of computational hardware, the size of the model, and the prompt length. In our case, the prompt is the BSM time window. Therefore, we focused our analysis on the impact of the window size on detection performance.
[…]
Despite the latency limitation, MistralBSM is designed to handle variable prompt lengths without requiring architectural modifications or additional preprocessing, ensuring flexibility in adapting the window size. As a consequence, the window size can be adjusted based on available computational resources to optimize performance without significant loss in accuracy.